A elevator car whose floor to ceiling distance is 2.7 m starts acending with a constant accl^n of 1.2 m/s^2 , 2 seconds after the start a bolt falls from the ceiling of the car . What is the free fall time of the bolt?

Grade:12th pass

1 Answers

Soumya Ranjan Mohanty
6 years ago
Dear student,
given dist between floor and ceiling=2.7 m
acceleration of elevator=1.2 ms^-2
relative acc of bolt with respect to elevator=g+a=9.8+1.2=11 ms^-2
relative velocity of bolt=0
so ,S=ut+1/2at^2
S=2.7 m,a=11 ,s^-2, u=0
=>2.7=0+ 1/2*11*t^2
solving, t= 0.7 seconds.
So the free fall time of the bolt=0.7 sec.
Thanks,
Soumya Ranjan Mohanty.